# Embodied Crowd Counting

### v1

@@ 4.1 Settings
where $M$ is the quantity of testing environments, $\hat{y}_{i}$ and $y_{i}$ are the estimated count and the ground truth count in an environment, respectively.
The sum of Euclidean distance between adjacent navigation points along the agent traveling path is used to evaluate the travel distance TD, which is defined as:
${\rm{TD}}=\sum\limits_{i=1}^{n-1}{\left\|{{{\bf{x}}_{i+1}}-{{\bf{x}}_{i}}}\right\|},$ (9)
where $n$ is the quantity of navigation points in one episode, ${{{\bf{x}}_{i}}}$ and ${{{\bf{x}}_{i+1}}}$ are the coordinates of the two adjacent navigation points, respectively.
Implementation Details. MLLM used in ATE is GPT-4V [46]. The path planning method is A* algorithm [37]. Altitude is $80m$ for HAE and $10m$ for LAE. The crowd density estimator in ATE is Generalized Loss (GL) [41], and the detection model in FDC is Grounding DINO (GD) [34]. For hyper parameters, navigation vector deg $\zeta$ is 15 °, density threshold $\kappa$ is 0.7, navigation point range $\eta$ is 8 m, and cluster size $\epsilon$ is 40.
Baselines. We compare ZECC with existing exploration method Frontier-based exploration (FBE) [39], as well as SOTA ZSON methods, 1) CoW [8], 2) OpenFMNav [22]. These methods use advanced foundation models to conduct decision making during navigation. To apply the baselines in ECC, the exploration step limitation for ZSON methods is removed, and they are equipped with GD or GL to conduct crowd counting using the projection method in FDC.
4.2 Results
Overall Performance. We report overall performance in Table 1. ZECC offers the optimal balance between counting performance and cost. In contrast, FBE delivers the best TD but lacks the ability to perform a target detection function, which limits its capacity for fine observation. This incomplete observation negatively impacts its overall performance. ZSON methods, such as Cow and OpenFMNav come with perception modules; however, they do not select a 3D fine observation point as NLBN, which means they cannot fully observe each individual in the crowd. This limitation leads to individuals being obscured by overlapping crowds, especially when the density is high. In some instances, these methods only identify single individuals, which reduces their efficiency. In contrast, ZECC improves counting efficiency by planning Z-axis exploration using ATE and employing NLBN for detailed detection.
For agents utilizing counting models for crowd counting, the performance tends to be slightly lower than that of agents using detection models. This is due to the increased sparsity of crowd observations as the agents get closer to the crowd. Crowd counting models are mainly trained on images depicting dense crowds, which results in limited generalization capability for scenarios with sparser crowds. Additionally, current ZSON methods do not consider finding a close position for crowd observation, which further degrades performance.
Method MAPE (%) TD (m)
FBE [39] + GL [41] 55.92 2227.71
FBE [39] + GD [34] 53.85 2227.71
CoW [8] + GL [41] 51.19 3271.07
CoW [8] + GD [34] 46.70 3271.07
OpenFMNav [22] + GL [41] 59.71 4936.79
OpenFMNav [22] + GD [34] 46.84 4936.79
ZECC 18.91 3804.63
Table 1: Overall performance of the methods. ZECC achieves best trade-off between MAPE and TD.
Figure 5: Performance and cost of ZECC and the baselines under different crowd density levels. L1-L5 refers to increasing density level. The figure demonstrate that ZECC achieves a balance between performance and exploration cost.
@@ 4.2 Results
Figure 6: The effect of four hyperparameters used in our model, including GMM cluster size, navigation vector degree, navigation point range and density map threshold.
We analyze various environments with different crowd density levels and visualize the average performance and cost of the methods using the GD detector in these scenarios. The results are presented in Figure 5. Among the methods evaluated, ZECC demonstrates the best average performance across different density levels. As crowd density increases, the MLLM in ATE becomes more effective at identifying high-density crowds, resulting in improved performance. In contrast, other methods struggle to manage high-density crowds, leading to limited performance.
In terms of cost, ZECC’s navigation points are influenced by crowd distribution and density. As the density level increases, the cost also rises. Although ZECC falls short of achieving the lowest cost in the last two density levels, its costs are still comparable to the baseline while ensuring effective counting performance. In contrast, other methods do not actively adjust navigation points based on crowd distribution, which results in their costs remaining relatively stable and limits their effectiveness in low-density scenes.
MAPE (%) TD (m)
ZECC w/o ATE 17.46 4633.67
ZECC w/o HAE 17.46 4633.67
ZECC w/o LAE 88.08 1738.84
ZECC 18.91 3804.63
Table 2: The results of ablation study for components in ATE. w/o ATE refers to using FBE + NLBN results for crowd counting. w/o HAE refers to fix the agent’s altitude to LAE. w/o LAE refers to fix the agent’s altitude to HAE. ZECC improves approximately 17% cost by reducing approximately 8% performance, achieving a better trade-off.
MAPE (%) Successful rate (%)
ZECC w/o NLBN 65.19 100.00
ZECC w/o NL 98.44 8.33
ZECC w/o VPS 92.55 100.00
ZECC w/o ATE-VPS 99.49 1.45
ZECC 18.91 100.00
Table 3: The results of ablation study for components in NLBN. w/o NLBN refers to using ATE results for crowd counting. w/o NL refers to not using normal line (NL) to calculate navigation points but use the cluster centers as navigation points. w/o VPS refers to not using view point selection (VPS) but select a point along the normal vector with $\eta$. w/o ATE-VPS refers to not using the ATE-view-vector-based view point selection (ATE-VPS), but randomly select a navigation vector from the potential navigation vectors. Successful rate indicates the ratio of the reachable navigation point reported by the A* path planning algorithm.
4.3 Ablation Study
ATE. We conducted ablation studies by removing specific components from the ATE. The results are presented in Table 2. The configuration without ATE (w/o ATE) achieves the best performance because it explores the environments greedily and utilizes NLBN for precise detection, resulting in adequate perception. However, this approach leads to a larger temporal delay (TD). In contrast, ZECC achieves a better balance by conducting both HAE and LAE simultaneously. w/o HAE uses the same setting as w/o ATE. w/o LAE fails for it maintains at a high altitude and does not conduct close range detection, making it hard to capture targets.
NLBN. We then conducted ablation studies on NLBN and the results are presented in Table 3. The findings indicate that omitting any component of NLBN results in a significant drop in either performance or success rates. Without NLBN, the absence of optimized viewpoints causes ZECC to revert to a ZSON method. When both NL and FBE-VPS are removed, most navigation points end up being located on obstacles, leading to a scarcity of reachable navigation points and incomplete exploration. Additionally, without VPS, the navigation points are positioned directly above the targets, causing the targets to fall out of the field of view, which results in corrupted observations.
4.4 Hyper parameter Study
We then examine the effects of parameters in the method, including GMM cluster size, navigation vector degree, navigation point range, and density map threshold. A gym-like scene featuring a densely packed crowd is utilized to test these parameters. The results are illustrated in Figure 6.
GMM cluster size. As the size of the GMM cluster increases, performance tends to deteriorate. When the GMM cluster size is set too low, the method generates sparse navigation points, which prevents the agent from effectively exploring crowded areas, resulting in poor counting performance. Although a smaller GMM cluster size can initially yield high performance, the results become worse when the cluster sizes reach 30 and 40, respectively. Specifically, the cost for a cluster size of 30 is 4065.02 meters, while for a size of 40, it rises to 5125.44 meters. Opting for an extreme GMM cluster size may significantly increase costs without providing substantial improvements in performance. Therefore, it is crucial to carefully determine the GMM cluster size to find a balance between cost and performance.
Navigation vector degree. Performance is optimal at 20°, while it declines at other angles. This degradation occurs because the agent has a limited field of view; it cannot effectively observe targets outside this range. This highlights the importance of our navigation point generation method.
Navigation point range. The performance is optimal at a distance of 8 meters, while performance declines at both shorter and longer ranges. A similar trend is observed with the navigation vector degree. At extremely close range, the agent’s field of view is restricted, and at long range, the agent is unable to gather detailed observations, which negatively impacts the subsequent target detection phase. Therefore, selecting the right viewpoint and generating appropriate navigation points are crucial for effective scene exploration algorithms.
Density map threshold. The results indicate that similar to GMM cluster size, performance declines as the threshold increases. A lower density map threshold results in a coarser estimation of the target zone, which in turn generates more navigation points. This allows for more detailed exploration and observation, leading to improved counting performance. Yet cost rises. For instance, when the density map threshold is 0.7, the travel distance is 4720.02 meters, and it increase to 7272.02 meters when the threshold is 0.4. This is a significant degradation in cost, yet improvement in performance is minor. This emphasizes the importance of parameter selection.
5 Conclusion
In this study, Embodied Crowd Counting (ECC) is proposed, a task which enables interactive crowd counting. A simulator, the Embodied Crowd Counting Dataset (ECCD) is developed to enable related research for ECC. This dataset includes 60 diverse virtual environments with crowd density modeled by a prior probability distribution, approximating the reality. A method, Zero-shot Embodied Counting (ZECC), is proposed to verify this task. This is an active agent which can explore unknown environments without additional assistance. Active Top-down Exploration (ATE) is proposed to utilize Z-axis moving ability for exploration planning. This module is equipped with MLLM to active high altitude exploration (HAE) or low altitude exploration (LAE), balancing crowd counting performance and exploration cost. Normal-line based Navigation (NLBN) is proposed to select optimized navigation point for crowd observation. This module generates a navigation point from the top-down view and maintains an angle, alleviating overlap of the crowd. Simultaneously, the estimate navigation points enable obstacle avoidance and ensure that the crowd is in FOV. Experiment results show that ZECC achieves a balance between performance and cost compared to recent navigation agents. Future research will focus on more efficient and scalable agents to enhance performance and adaptability.

### v2

@@ C.1 ATE
[…]
8:while not $done1$ do
9:  $O_{t},p_{t}\leftarrow\text{getState}()$;
10:  $s_{t}\leftarrow\text{MLLM}\left({{O_{t}};I}\right)$;
11:  $M_{H}\leftarrow\text{updateMap}\left({{O_{t}};p_{t}}\right)$;
12:  if $s_{t}>0.5$ then
13:   $\text{toLAE}()$;
14:   $done2\leftarrow False$;
15:   while not $done2$ do
16:     $O_{f},p_{f}\leftarrow\text{getState}()$;
17:     $d_{f}\leftarrow\text{P}\left({\text{G}\left({{O_{f}}}\right),{p_{f}}}\right)$;
18:     $M_{L}\leftarrow\text{updateMap}\left({{O_{f}};p_{f}}\right)$;
19:     $D\leftarrow\text{updateDistribution}(D,d_{f})$;
20:     $done2\leftarrow\text{toUnexplored}(M_{L})$;
21:   end while
22:   toHAE();
23:  end if
24:  $done1\leftarrow\text{toUnexplored}(M_{H})$;
25:  $t\leftarrow t+1$;
26:end while
27:return $D$
C.2 NLBN
The method used to estimate the normal line of the crowd cluster plane is the Open3D package. The global crowd distribution is first divided into subareas using GMM cluster. Then, on each cluster, Open3D is used to estimate the normal line for each cluster. The normal lines are aligned with the vector (0,0,1). Then, the normal line at the cluster center is selected as the normal line of the cluster. Based on this representation, the navigation points can be obtained. The pseudo-code is shown in Algorithm 2.
Algorithm 2 Pseudo-Code of NLBN
@@ C.2 NLBN
1:Cluster size $\epsilon$, Navigation vector degree $\zeta$, Navigation point range $\eta$, Density map threshold $\kappa$, Global distribution $D$
2:
[…]
6:for $i=1$ to NN do
7:  ${\bf{d}}_{i}^{{\rm{cluster}}}\leftarrow\text{FitPlane}({\bf{x}}_{i}^{{\rm{cluster}}})$;
8:  ${\bf{x}}_{i}^{{\rm{ATE}}}\leftarrow\text{getViewpoint}(D)$;
9:  Compute $\{\mathbf{d}_{i1}^{\text{view}},...,\mathbf{d}_{im}^{\text{view}}\}_{i}$ using Eq. (4);
10:  Compute ${\bf{d}}_{i}^{{\rm{ATE}}}$ using Eq. (5);
11:  Compute ${\bf{d}}_{i}^{{\rm{view}}}$ using Eq. (6);
12:  Compute ${\bf{x}}_{i}^{{\rm{view}}}$ using Eq. (7);
13:end for
14:while not donedone do
15:  $done\leftarrow\text{toNextNaviPoint}(\{{\bf{x}}_{i}^{{\rm{view}}},...,{\bf{x}}_{N}^{{\rm{view}}}\})$
16:end while
C.3 FDC
During FDC, the center of each detection box can be projected to the global crowd distribution by using the depth information. To re-identify targets, the space is divided into 3D voxels with size of 0.25 m. All the target with in a voxel is regared as the same target. This configuration is set for ZECC and all comparison methods.
Figure 9: Prompt template used in ATE.
Appendix D Occlusion analysis on ZECC
In this section, we further test the ability of ZECC to alleviate occlusion in the complex environments. For each ground truth person, we first select the navigation point which has the minimal Euclidean distance to it, and then test whether occlusion occurs between the navigation point and the person. This is implemented by first projecting the global point cloud to the vector from navigation point to the person. Then, if the minimal distance between the global point cloud and the projected points is lower than a threshold, the person is obstructed. To formulate this, the unit vector from a person to its nearest navigation point is calculated as:
${\bf{v}}_{i}^{{\rm{navi}}}=\frac{{{\bf{P}}_{i}^{{\rm{navi}}}-{\bf{P}}_{i}^{{\rm{crowd}}}}}{{\left\|{{\bf{P}}_{i}^{{\rm{navi}}}-{\bf{P}}_{i}^{{\rm{crowd}}}}\right\|}},$ (8)
where ${\bf{P}}_{i}^{{\rm{navi}}}$} is coordinate of the navigation point and ${\bf{P}}_{i}^{{\rm{crowd}}}$} is the coordinate of the person. The vector from a global point cloud to the person is:
${\bf{v}}_{ij}^{{\rm{global}}}={\bf{P}}_{i}^{{\rm{navi}}}-{\bf{P}}_{j}^{{\rm{global}}},$ (9)
where ${\bf{P}}_{j}^{{\rm{global}}}$ is a random coordinate from the global point cloud. The projected length is:
${l_{ij}}={\bf{v}}_{ij}^{{\rm{global}}}\cdot{\bf{v}}_{i}^{{\rm{navi}}},$ (10)
and the projected point is:
${\bf{P}}_{j}^{{\rm{proj}}}={\bf{P}}_{i}^{{\rm{crowd}}}+{l_{ij}}*{\bf{v}}_{i}^{{\rm{navi}}}\text{ s.t. }{l_{ij}}\geq 0\wedge{l_{ij}}\leq\left\|{{\bf{v}}_{ij}^{{\rm{global}}}}\right\|.$ (11)
The condition of determining the person is obstructed is:
$\mathop{{\rm{min}}}\limits_{j}\left\|{{\bf{P}}_{j}^{proj}-{\bf{P}}_{j}^{global}}\right\|\leq\lambda.$ (12)
In the experiment, $\lambda$ is set to 0.5 m. We compare ZECC with OpenFMNav and MVC-30. The ratio between the number of obstructed person and the number of ground truth person of the three methods are shown in Table 7. It shows that ZECC suffers from the less occlusion, which benefits the crowd counting results.
Appendix E Failure case study
ZECC fails when MLLM makes wrong decision. This is mainly due to occlusion by buildings. Figure 10 shows an example of MLLM planning failure, where the agent does not choose to conduct LAE at a crowd area.
Figure 10: A case of ZECC failure. (a) Agent trajectory. (b) Top-down RGB. (c) MLLM reasoning. Blue box: ground truth crowd. Because of occlusion, MLLM failed to perceive the crowd during HAE.
Appendix F Quantitative results in real environment
To test ZECC in the real environment, we use two outdoor crowd scene to verify the proposed method. The drone model used is the Taobotics Q300. The captured images are transmitted back to a ground server using a remote communication protocol and fed into VGGT [49] to estimate relative point clouds and poses. The proposed NLBN is then employed to calculate relative navigation points. The drone’s absolute geographic coordinates and poses are recorded in real-time by its GPS. By leveraging the relationships between absolute and relative geographic coordinates and poses, the navigation points calculated by NLBN are mapped to absolute navigation points. After the ground server controls the drone to fly to these absolute navigation points, it captures images of the crowd. We utilize Grounding DINO and the crowd counting model Generalized Loss to perform crowd detection or counting on images captured from both distant and close-up perspectives, respectively. The ground truth for crowd annotations is manually labeled. The visualization of the drone observation before and after adjusting its position to the NLBN navigation point is shown in Figure 11. The quantitative results are shown in Table 7. Before applying NLBN, the drone can not obtain a clear view of the crowds in the target area. The GL estimation result is fuzzy, and due to crowd overlap, people in the back rows are obstructed by those in the front rows. This results in the corrupted detection of GD. By adjusting to the NLBN navigation point, the drone is able to observe the target area from a top-down perspective, alleviating occlusion of people in back rows and improving counting performance significantly. This demonstrates how NLBN addresses occlusion.
Table 6: Obstructed person ratio.
Method Ratio (%)
OpenFMNav 42.81
MVC-30 21.77
ZECC 4.52
Table 7: Performance of real scenarios.
Methods MAPE (%)
Scenario 1 Scenario 2
w/o NLBN + GL 41.41 57.76
w/o NLBN + GD 73.74 78.44
NLBN + GL 22.50 29.31
NLBN + GD 15.17 19.83
Figure 11: Quantitative results in real environment. (a), (d) Output of GL. The marked areas indicate the target zone. (b), (e) Point cloud of the scene. Red lines: candidate view vector. Green lines: the drone view vector. Orange lines: selected view vector. (c), (f) GD detection results. Red boxes: detected crowds. Green boxes: ground truth crowds. First row: before applying NLBN. Second row: after applying NLBN. (a), (b) and (c) are Scenario 1 and (d), (e) and (f) are Scenario 2. By adjusting the camera position to the NLBN navigation point, the target detection result is improved.